OpenStudy (anonymous):

x-3/y+4 = 12, is the answer x= 12y+51????

OpenStudy (anonymous):

you multiply y+4 to the other side which gives you x-3 = 12y+48 then add 3 to the other side x=12y+51

OpenStudy (anonymous):

x-3=12y+48 x=12y+51

OpenStudy (anonymous):

Thanks, just making sure.